OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. B(0, 0), C(4, -4)

OpenStudy (anonymous):

we need to find the slope using the slope formula : slope(m) = (y2 - y1) / (x2 - x1) (0,0) x1 = 0 and y1 = 0 (4,-4) x2 = 4 and y2 = -4 now we sub slope(m) = (-4-0) / (4 - 0) slope(m) = -4/4 = -1 so now we have our slope(m) = -1 and we have our x and y points. We will use y = mx + b to find b. y = mx + b m = -1 (4,-4) x = 4 and y = -4 (you can use either set of points, I just picked this set) now we sub -4 = 4(-1) + b -4 = -4 + b 0 = b our equation is : y = -1x + 0 (we will turn it into standard form now) x + y = 0 ===> standard form

OpenStudy (anonymous):

any questions ?

OpenStudy (anonymous):

what if it wants x first?

OpenStudy (anonymous):

Indicate the equation of the given line in standard form. The line through (2, -1) and parallel to a line with slope of 3/4

OpenStudy (anonymous):

a parallel line will have the same slope. So the slope we will use is 3/4. y = mx + b slope(m) = 3/4 (2,-1) x = 2 and y = -1 now we sub -1 = 3/4(2) + b -1 = 6/4 + b -1 - 6/4 = b -4/4 - 6/4 = b -10/4 = b - 5/2 = b so our equation is : y = 3/4x - 5/2 (multiply by the LCD, which is 4) 4y = 3x - 10 (subtract 3x from both sides) -3x + 4y = - 10 (multiply by -1 to make x positive) 3x - 4y = 10 ===>standard form

OpenStudy (anonymous):

what do you mean by " what if it wants x first " ?